You are the brand manager for ‘Head and Ear’ shampoo and you have 20% market share (in units). Your unit contribution from this shampoo is $2.5. You wanted to launch a new shampoo called ‘Lovely Hair’ and you conduct an ASSESSOR pretest market study. The pretest market predicted a market share of 10% (in units) for Lovely Hair and also predicted that your ‘Head and Ear’ shampoo market share post launch will be 15% (in units). The unit contribution of ‘Lovely Hair’ shampoo is $5. Assume that the total shampoo category sales is X units per year and it remains unchanged after introduction of ‘lovely hair’. If you launch the ‘Lovely Hair’ shampoo, calculate the percentage change in the total contribution for your company in the shampoo category?
Answer #1

Let the total shampoo category sales per year be 100.

That is, X=100

As per the question, the value of X remains unchanged.

When 'Lovely Hair' shampoo is introduced, the market share of this category is predicted to be 10%.

That is, the number of units sold for 'Lovely Hair' shampoo is = 10%*100 = 10

Also, the market share of 'Head and Ear' shampoo post launch would be 15%.

That is, the number of units sold for 'Head and Ear' shampoo is = 15% * 100 = 15

Total units sold = 15+10 = 25

Total contribution from these categories of shampoo = (15*2.5)+(10*5) = $87.5

Previously (before the launch of 'Lovely Hair' shampoo), the total units of 'Head and Ear' shampoo sold = 20%*100 =20

So, total contribution of this shampoo=20*2.5= $50

Percentage change in the total contribution= (87.5-50)/87.5 * 100 = 42.86%

Answer: The percentage change in the total contribution is 42.86%.